# Nonuniqueness of Leray–Hopf solutions to the unforced incompressible 3D Navier–Stokes Equation

### v1

@@ 5 Numerical Construction of Approximate Solutions
In this section, we detail the numerical procedures employed to solve the equation (1.12). In prior research [25], boundary conditions of the form:
@@ 5.1 Gradient descent for eigenvalue
    Spectral bound. (to suppress the most negative eigenvalue in the far field):
@@ 5.2 Axisymmetry, parity and boundary conditions
We consider the axisymmetric case throughout the computations and denote the cylindrical coordinates as $(\rho,z,\phi)$. We also impose symmetry conditions in $z$: even for the profile problem $(U,P)$ and odd for the eigenvalue problem $(v,q)$. In Euclidean coordinates, a vector field $(u_{x},u_{y},u_{z})$ is even in $z$ if
@@
In practice, for the construction of accurate approximate solutions in this section, we adopt spherical coordinates $(r,\theta,\phi)$. Due to the axisymmetry and parity, the computational domain reduces to $r\in\mathbb{R}_{+}$ and $\theta\in\left[0,\frac{\pi}{2}\right]$. Since spherical coordinates $(r,\theta)$ correspond to polar representation of cylindrical coordinates $(\rho,z)$, boundary conditions in spherical coordinates are thus given by:
@@
    At $\theta=0$, which corresponds to $\rho=0$, we have $e_{r}=e_{z}$ and $e_{\theta}=e_{\rho}$. To ensure $\Delta U,\Delta v\in L^{2}$ and smoothness of the pressure, we impose Dirichlet zero condition on $\rho$ and $\phi$ component of the velocity field and Neumann zero condition on $z$ component and the pressure. This implies $\widetilde{U}_{\theta}(r,0)=\widetilde{U}_{\phi}(r,0)=0$ and $\partial_{\theta}\widetilde{U}_{r}(r,0)=\partial_{\theta}\widetilde{P}(r,0)=0$. Likewise, $\widetilde{v}_{\theta}(r,0)=\widetilde{v}_{\phi}(r,0)=0$ and $\partial_{\theta}\widetilde{v}_{r}(r,0)=\partial_{\theta}\widetilde{q}(r,0)=0$.
@@
    At $r=0$, which corresponds to the origin in cylindrical coordinates, the even symmetry in $z$ and Dirichlet zero condition in $\rho$ directly lead to $\widetilde{U}(0,\theta)=(0,0,0).$ Similarly, $\partial_{r}\widetilde{P}(0,\theta)=0$ and $\widetilde{q}(0,\theta)=0$. Finally, since at the origin in cylindrical coordinates, $\widetilde{v}$ must align with the direction $e_{z}$, we know that $\widetilde{v}$ takes the following form for a constant $C$:
@@ 5.3 Spherical coordinates and finite element method
In this section, we detail our setting for finite element approximation. The incompressible Navier–Stokes equations in spherical coordinates read:
@@
where $\nabla_{\text{sph}}$, $\text{div}_{\text{sph}}$, and $\Delta_{\text{sph}}$ are the gradient, divergence, and Laplacian operators in spherical coordinates, respectively (see Appendix D for explicit expressions). The weak form of (5.6) is:
@@ 5.5 Extension to whole space
Numerical instabilities occasionally lead to spurious eigenvalues with highly oscillatory eigenvectors, which do not correspond to actual eigenmodes of the operator. However, despite these numerical artifacts, the method still enables us to approximate the genuine eigenvectors accurately.
@@ 5.6 Construction of divergence-free basis
To exactly enforce the divergence-free condition, we construct a divergence-free basis. We first use vector spherical harmonics (see Appendix E for introduction) to expand the vector $\widetilde{U}$:
@@
We note that when $l$ is even, the vector field $(Y_{l},X_{l},0)$ is even with respect to $z$ while $(0,0,Z_{l})$ is odd. When $l$ is odd, they have the opposite parities. Therefore, for the even $\widetilde{U}$, we choose even $l$ for $Y_{l},X_{l}$ and odd $l$ for $Z_{l}$. For the odd $\widetilde{v}$, it is the other way around. Therefore, we modify our expansion to:
@@
$\frac{1}{2}\sin 2\beta\frac{\mathrm{d}}{\mathrm{d}\beta}\widehat{U}_{l}^{Y}+\frac{1}{2}(3+\cos 2\beta)\widehat{U}_{l}^{Y}-\mu_{2l}\widehat{U}_{l}^{X}=0.$ (5.11)
In practice, we truncate the series up to $L$ terms. Therefore, the divergence free condition splits into $L$ equations relating the radial coefficient functions $\widehat{U}_{l}^{Y}$ and $\widehat{U}_{l}^{X}$.
In order to expand the coefficient functions into trigonometric series, we first examine the boundary behavior. Specifically, we recall the boundary conditions in Section Section 5.2 as $\widehat{U}(\beta=0,\theta)=(0,0,0)$ at the origin. The far field behavior (1.13) in Assumption 1 implies that
@@
$\widehat{U}_{l}^{Y}(\beta)=\sum_{m=1}^{M}\mathrm{B}^{U,Y}_{m,l}\sin((2m-1)\beta),\quad\widehat{U}_{l}^{X}(\beta)=\sum_{m=1}^{M+1}\mathrm{B}^{U,X}_{m,l}\sin((2m-1)\beta).$ (5.12)
@@
with boundary conditions: $\mathrm{B}^{U,Y}_{0,l}=\mathrm{B}^{U,Y}_{M+1,l}=\mathrm{B}^{U,Y}_{M+2,l}=0$.
@@
with boundary conditions: $\mathrm{B}^{v,Y}_{0,l}=\mathrm{B}^{v,Y}_{M+1,l}=\mathrm{B}^{v,Y}_{M+2,l}=0$.
These algebraic constraints ensure strict divergence-freeness of $\overline{U}$ and $\overline{v}$. In the next subsection, we demonstrate how to interpolate our solutions onto this basis.
@@ 5.7 Divergence-free interpolation
In the preceding sections, we obtained numerical approximations for the profile $\overline{U}$ and the eigenvector $\overline{v}$ using the finite element method. To ensure strict adherence to the divergence-free condition, we now interpolate these solutions onto the carefully structured divergence-free basis developed in the previous subsection.
We use $\overline{U}_{\theta}$ as an example. We select a computational mesh defined by $\beta_{j}=\frac{j\pi}{2M},\theta_{k}=\frac{k\pi}{2L}$, and evaluate $\overline{U}_{\theta}$ at each grid point $(\beta_{j},\theta_{k})$. We apply discrete Fourier transforms in $\theta$ to obtain expansions:
$\overline{U}_{\theta}(\beta_{j},\theta)=\sum_{l=1}^{L}\mathcal{F}_{\theta}\overline{U}_{\theta}^{l}(\beta_{j})\sin(2l\theta),\quad j=1,\ldots,M.$
where $\mathcal{F}_{\theta}$ denotes the discrete Fourier transform with respect to $\theta$. We then convert the obtained Fourier expansions to the eigenfunction basis $X_{2l}$. Therefore, we get the value of $\widehat{U}_{l}^{X}$ at each $\beta_{j}$. Applying another discrete Fourier transform in $\beta$, we derive the coefficients $\mathrm{B}^{U,X}_{m,l}$. This process is similarly conducted for other components $\overline{U}_{r}$, $\overline{v}_{r}$ and $\overline{v}_{\theta}$, with their respective boundary conditions. For components $\overline{U}_{\phi}$, $\overline{P}$, $\overline{v}_{\phi}$ and $\overline{q}$, the linear transformation is not required since we do not need to represent them in the Legendre basis to enforce divergence-free conditions.
We explicitly enforce the divergence-free conditions by setting coefficients $\mathrm{B}^{U,X}_{m,l}$ and $\mathrm{B}^{v,X}_{m,l}$ according to the derived algebraic constraints (5.13) and (5.14). In practice, we observe only small deviations from the original values.
@@ 5.8 Summary
The numerical algorithm is summarized as follows:
@@
     Update solution and boundary condition: $U_{n},A_{n}\to U_{n+1}^{\ast},A_{n+1}$; // Note $U_{n+1}^{\ast}$ is not yet on the solution manifold; it will be corrected in the next iteration of the loop.
@@
Solve equation (5.9) using Newton’s method: $\overline{U}_{n}\to\overline{U}_{n}^{\operatorname{cv}}$;
Compute eigenpair $(\overline{\lambda}_{n},\overline{v}_{n}^{\operatorname{cv}})$;
Interpolate the solutions onto a divergence-free basis: $\overline{U}_{n}^{\operatorname{cv}},\overline{v}_{n}^{\operatorname{cv}}\to\overline{U}_{n}^{\operatorname{df}},\overline{v}_{n}^{\operatorname{df}}$;
Initialize iteration index $n=0$;
@@
while $\|R_{\operatorname{sph}}^{U}(\overline{U}_{n}^{\operatorname{df}},\overline{P}^{\operatorname{df}}_{n})\|>\varepsilon_{\operatorname{res}}$ or $\|R_{\operatorname{sph}}^{v}(\overline{v}_{n}^{\operatorname{df}},\overline{q}^{\operatorname{df}}_{n})\|>\varepsilon_{\operatorname{res}}$ do
   Solve (5.17) to obtain $(\delta\overline{U}_{n}^{\operatorname{cv}},\delta\overline{P}^{\operatorname{cv}}_{n})$ and an analogous problem for $(\delta\overline{v}_{n}^{\operatorname{cv}},\delta\overline{q}^{\operatorname{cv}}_{n})$;
  Update $\overline{U}_{n+1}^{\operatorname{cv}},\overline{P}^{\operatorname{cv}}_{n+1},\overline{v}_{n+1}^{\operatorname{cv}},\overline{q}^{\operatorname{cv}}_{n+1}\leftarrow\overline{U}_{n}^{\operatorname{cv}}+\delta\overline{U}_{n}^{\operatorname{cv}},\overline{P}^{\operatorname{cv}}_{n}+\delta\overline{P}_{n}^{\operatorname{cv}},\overline{v}_{n}^{\operatorname{cv}}+\delta\overline{v}_{n}^{\operatorname{cv}},\overline{q}^{\operatorname{cv}}_{n}+\delta\overline{q}_{n}^{\operatorname{cv}}$;
  Update $\overline{\lambda}_{n+1}$ based on orthogonality by Assumption 1;
  Increment the iteration index $n\leftarrow n+1$;
@@
return Profile $\overline{U}^{\text{cv}}$ and eigen pair $(\overline{\lambda},\overline{v}^{\text{cv}})$;

### v2

@@ 5 Numerical Construction of Approximate Solutions
In this section, we detail the numerical procedures used to solve the equation (1.12). In prior research [27], boundary conditions of the form:
@@ 5.1 Gradient descent for eigenvalue
    Spectral bound (to suppress the most negative eigenvalue in the far field):
@@ 5.2 Axisymmetry, parity and boundary conditions
We consider the axisymmetric case throughout the computations and denote the cylindrical coordinates as $(\rho,z,\phi)$. We also impose symmetry conditions in $z$: even for the profile problem $(U,P)$ and odd for the eigenvalue problem $(v,q)$. In the Euclidean coordinates, a vector field $(u_{x},u_{y},u_{z})$ is even in $z$ if
@@
In practice, for the construction of accurate approximate solutions in this section, we adopt spherical coordinates $(r,\theta,\phi)$. Due to the axisymmetry and parity, the computational domain reduces to $r\in\mathbb{R}_{+}$ and $\theta\in\left[0,\frac{\pi}{2}\right]$. Since the spherical coordinates $(r,\theta)$ correspond to a polar representation of the cylindrical coordinates $(\rho,z)$, boundary conditions in the spherical coordinates are thus given by:
@@
    At $\theta=0$, which corresponds to $\rho=0$, we have $e_{r}=e_{z}$ and $e_{\theta}=e_{\rho}$. To ensure $\Delta U,\Delta v\in L^{2}$ and smoothness of the pressure, we impose a Dirichlet zero condition on the $\rho$ and $\phi$ components of the velocity field and a Neumann zero condition on the $z$ component and the pressure. This implies $\widetilde{U}_{\theta}(r,0)=\widetilde{U}_{\phi}(r,0)=0$ and $\partial_{\theta}\widetilde{U}_{r}(r,0)=\partial_{\theta}\widetilde{P}(r,0)=0$. Likewise, $\widetilde{v}_{\theta}(r,0)=\widetilde{v}_{\phi}(r,0)=0$ and $\partial_{\theta}\widetilde{v}_{r}(r,0)=\partial_{\theta}\widetilde{q}(r,0)=0$.
@@
    At $r=0$, which corresponds to the origin in the cylindrical coordinates, the even symmetry in $z$ and the Dirichlet zero condition in $\rho$ directly lead to $\widetilde{U}(0,\theta)=(0,0,0).$ Similarly, we impose $\partial_{r}\widetilde{P}(0,\theta)=0$ and $\widetilde{q}(0,\theta)=0$. Finally, since at the origin in cylindrical coordinates, $\widetilde{v}$ must align with the direction $e_{z}$, we know that $\widetilde{v}$ takes the following form for a constant $C$:
@@ 5.3 Spherical coordinates and finite element method
In this section, we detail our setting for finite element approximation. The incompressible Navier–Stokes equations in the spherical coordinates read:
@@
where $\nabla_{\text{sph}}$, $\text{div}_{\text{sph}}$, and $\Delta_{\text{sph}}$ are the gradient, divergence, and Laplacian operators in the spherical coordinates, respectively (see Appendix D for explicit expressions). The weak form of (5.6) is:
@@ 5.5 Extension to whole space
Numerical instabilities occasionally lead to spurious eigenvalues with highly oscillatory eigenvectors, which do not correspond to the actual eigenmodes of the operator. However, despite these numerical artifacts, the method still enables us to approximate the genuine eigenvectors accurately.
@@ 5.6 Construction of divergence-free basis
To exactly enforce the divergence-free condition, we construct a divergence-free basis. We first expand the vector $\widetilde{U}$ using axisymmetric vector spherical harmonics $X_{l},Y_{l},Z_{l}$, defined later in (5.10):
@@
Here, for a scalar function $p$, we can expand $p$ with the scalar spherical harmonics:
$p(r,\omega)=\sum_{l=0}^{+\infty}\sum_{m=-l}^{l}p_{lm}(r)S_{lm}(\omega),\quad r\geqslant 0,\quad\omega\in\mathbb{S}^{2},$
where
$Y_{lm}(\omega)=S_{lm}(\omega)\vec{e}_{r},\quad X_{lm}(\omega)=\nabla_{\mathbb{S}^{2}}S_{lm}(\omega),\quad Z_{lm}(\omega)=\vec{e}_{r}\times\nabla_{\mathbb{S}^{2}}S_{lm}(\omega).$
Since we only consider axisymmetric functions, we choose $m=0$ and $\omega$ reduces to $\theta$:
$Y_{l}(\theta)=S_{l}(\theta)\vec{e}_{r},\quad X_{l}(\theta)=S_{l}^{\prime}(\theta)\vec{e}_{\theta},\quad Z_{l}(\theta)=\vec{e}_{\phi}S_{l}^{\prime}(\theta),\quad S_{l}(\theta)=P_{l}(\cos\theta),$ (5.10)
We have the reduced expansions
We note that $X_{l},Y_{m},Z_{n}$ are mutually orthogonal for all indices $l,m,n$. Furthermore, they are normalized in the $L^{2}$ inner product:
The gradients of $p$ and $u$ have the following forms, by Appendix D and the eigenvector properties of the spherical harmonics and the Legendre polynomials:
$\nabla_{\operatorname{sph}}p$ $=$ $\sum_{l=0}^{+\infty}\left(p_{l}^{\prime}(r)Y_{l}(\theta)+\frac{1}{r}p_{l}(r)X_{l}(\theta)\right),$ (5.11)
$\Delta_{\operatorname{sph}}p$ $=$ $\sum_{l=0}^{+\infty}\left(\Delta_{r}p_{l}(r)-\frac{\mu_{l}}{r^{2}}p_{l}(r)\right)S_{l}(\theta),$ (5.12)
$\operatorname{div}_{\operatorname{sph}}u$ $=$ $\sum_{l=0}^{+\infty}\left(\frac{\mathrm{d}}{\mathrm{d}r}u_{l}^{Y}+\frac{2}{r}u_{l}^{Y}-\frac{\mu_{l}}{r}u_{l}^{X}\right)S_{l},$ (5.13)
$\Delta_{\operatorname{sph}}u$ $=$ $\sum_{l=0}^{+\infty}(\Delta_{r}u_{l}^{Y}Y_{l}+\Delta_{r}u_{l}^{X}X_{l}+\Delta_{r}u_{l}^{Z}Z_{l})$ (5.14)
where we define
We note that the superscripts $X,Y,Z$ refer to the spherical harmonics and have nothing to do with Cartesian coordinates. We note that when $l$ is even, the vector field $(Y_{l},X_{l},0)$ is even with respect to $z$ while $(0,0,Z_{l})$ is odd. When $l$ is odd, they have the opposite parities. Therefore, for the even $\widetilde{U}$, we choose even $l$ for $Y_{l},X_{l}$ and odd $l$ for $Z_{l}$. For the odd $\widetilde{v}$, it is the other way around. Therefore, we modify our expansion to:
@@
$\frac{1}{2}\sin 2\beta\frac{\mathrm{d}}{\mathrm{d}\beta}\widehat{U}_{l}^{Y}+\frac{1}{2}(3+\cos 2\beta)\widehat{U}_{l}^{Y}-\mu_{2l}\widehat{U}_{l}^{X}=0,\quad\forall l\geqslant 0.$ (5.16)
For each mode $l$, since both $\widehat{U}_{l}^{Y}$ terms are multiplied with $\sin 2\beta$ or $\cos 2\beta$, we need one more Fourier mode in $\beta$ for $\widehat{U}_{l}^{X}$ to make sure (5.16) holds. In practice, we truncate the series in (5.15) up to $L$ terms. Therefore, the divergence free condition splits into $L$ equations relating the $\beta$ coefficient functions $\widehat{U}_{l}^{Y}$ and $\widehat{U}_{l}^{X}$.
In order to expand the coefficient functions into trigonometric series, we first examine the boundary behavior. Specifically, we recall the boundary conditions in Section 5.2 as $\widehat{U}(\beta=0,\theta)=(0,0,0)$ at the origin. The far field behavior (1.13) in Assumption 1 implies that
@@
$\widehat{U}_{l}^{Y}(\beta)=\sum_{m=1}^{M}\mathrm{B}^{U,Y}_{m,l}\sin((2m-1)\beta),\quad\widehat{U}_{l}^{X}(\beta)=\sum_{m=1}^{M+1}\mathrm{B}^{U,X}_{m,l}\sin((2m-1)\beta),$ (5.17)
since
$\sin((2m-1)\beta=(-1)^{m-1}+O((\frac{\pi}{2}-\beta)^{2}).$
@@
with boundary conditions: $\mathrm{B}^{U,Y}_{0,l}=\mathrm{B}^{U,Y}_{M+1,l}=\mathrm{B}^{U,Y}_{M+2,l}=0$. Now we have solved the coefficients $\mathrm{B}^{U,X}$ from $\mathrm{B}^{U,Y}$ to enforce the divergence-free condition.
@@
with boundary conditions: $\mathrm{B}^{v,Y}_{0,l}=\mathrm{B}^{v,Y}_{M+1,l}=\mathrm{B}^{v,Y}_{M+2,l}=0$. These algebraic constraints ensure strict divergence-freeness of $\overline{U}$ and $\overline{v}$. From (5.18) and (5.19), we can construct a divergence-free basis from any trigonometric basis of $U_{r}$ and $U_{\phi}$. For example, we take the coefficients $\mathrm{B}^{U,Y},\mathrm{B}^{U,Z}$ as independent variables, and define $\mathrm{B}^{U,X}$ by the divergence-free constraint (5.18). This defines a linear reconstruction map from the space of free coefficients into the truncated vector-field space. Its image is precisely the truncated exactly divergence-free subspace. Therefore, by applying this reconstruction map to the standard coordinate basis in the free coefficient space, we obtain a basis of the truncated exactly divergence-free subspace. In the next subsection, we demonstrate how to interpolate our solutions onto this basis.
@@ 5.7 Divergence-free interpolation
In the preceding subsections, we obtained numerical approximations for the profile $\overline{U}$ and the eigenvector $\overline{v}$ using the finite element method. To ensure strict adherence to the divergence-free condition, we now interpolate these solutions onto the carefully structured divergence-free basis developed in the previous subsection.
We remark that $\overline{U}_{\theta}$, $\overline{U}_{r}$, $\overline{U}_{\phi}$ corresponds to the $X$, $Y$ and $Z$ family coefficients in the harmonic expansions respectively, and we enforce the divergence-free condition in a mode-by-mode fashion, to be precise, via the algebraic relations (5.18) and (5.19). We use $\overline{U}_{r}$ as an example. We select a computational mesh defined by $\beta_{j}=\frac{j\pi}{2M},\theta_{k}=\frac{k\pi}{2L}$, and evaluate $\overline{U}_{r}$ at each grid point $(\beta_{j},\theta_{k})$. We apply discrete Fourier transforms in $\theta$ to obtain expansions:
$\overline{U}_{r}(\beta_{j},\theta)=\sum_{l=1}^{L}\mathcal{F}_{\theta}\overline{U}_{r}^{l}(\beta_{j})\sin(2l\theta),\quad j=1,\ldots,M,$
where $\mathcal{F}_{\theta}$ denotes the discrete Fourier transform with respect to $\theta$. We then convert the resulting Fourier expansion in $\theta$ into the eigenfunction basis $Y_{2l}$, which amounts to a finite change of basis in the angular variable $\theta$. Therefore, we get the value of $\widehat{U}_{l}^{Y}$ at each $\beta_{j}$. Applying another discrete Fourier transform in $\beta$, we derive the coefficients $\mathrm{B}^{U,Y}_{m,l}$. Similarly we obtain $\mathrm{B}^{v,Y}_{m,l}$.We explicitly enforce the divergence-free conditions by setting coefficients $\mathrm{B}^{U,X}_{m,l}$ and $\mathrm{B}^{v,X}_{m,l}$ according to the derived algebraic constraints (5.18) and (5.19). In practice, we observe only small deviations from the original values.
@@ 5.8 Summary
For clarity, we distinguish four stages of the numerical approximation. We denote by $(\overline{U}_{n},\overline{P}_{n})$ the finite-element solution computed on the bounded computational domain using $n$-steps of Newton iterations. After applying the coordinate transformation described above, we obtain the whole-space candidate $(\overline{U}^{\mathrm{tr}},\overline{P}^{\mathrm{tr}})$. Interpolating this field into the truncated exactly divergence-free spectral representation yields $(\overline{U}^{\mathrm{df}},\overline{P}^{\mathrm{df}})$. Finally, we use $(\overline{U}^{(k)},\overline{P}^{(k)})$ for the iteratively refined approximations produced by the residual correction procedure at iteration $k$. The converged output of this residual correction procedure is denoted by $(\overline{U}^{\mathrm{cv}},\overline{P}^{\mathrm{cv}})$ and $(\overline{\lambda},\overline{v}^{\mathrm{cv}},\overline{q}^{\mathrm{cv}})$.
We summarize below the four-stage numerical pipeline, from the finite-element approximation on the bounded computational domain to the final corrected profile and associated eigenpair.
@@
     Update solution and boundary condition: $\overline{U}_{n},A_{n}\to\overline{U}_{n+1}^{\ast},A_{n+1}$; // Note $\overline{U}_{n+1}^{\ast}$ is not yet on the solution manifold and will be corrected in the next iteration of the loop.
@@
Solve equation (5.9) using Newton’s method: $\overline{U}_{n}\to\overline{U}^{\operatorname{tr}}$;
Compute eigenpair $(\overline{\lambda},\overline{v}^{\operatorname{tr}})$;
Interpolate the solutions onto a divergence-free basis: $\overline{U}^{\operatorname{tr}},\overline{v}^{\operatorname{tr}}\to\overline{U}^{\operatorname{df}},\overline{v}^{\operatorname{df}}$;
Initialize iteration index $k=0$, with initialization $\overline{U}^{(0)},\overline{v}^{(0)}=\overline{U}^{\operatorname{df}},\overline{v}^{\operatorname{df}}$;
@@
while $\|R_{\operatorname{sph}}^{U}(\overline{U}^{(k)},\overline{P}^{(k)})\|>\varepsilon_{\operatorname{res}}$ or $\|R_{\operatorname{sph}}^{v}(\overline{v}^{(k)},\overline{q}^{(k)})\|>\varepsilon_{\operatorname{res}}$ do
   Solve (5.22) to obtain $(\delta\overline{U}^{(k)},\delta\overline{P}^{(k)}$ and an analogous problem for $(\delta\overline{v}^{(k)},\delta\overline{q}^{(k)})$;
  Update $\overline{U}^{(k+1)},\overline{P}^{(k+1)},\overline{v}^{(k+1)},\overline{q}^{(k+1)}\leftarrow\overline{U}^{(k)}+\delta\overline{U}^{(k)},\overline{P}^{(k)}+\delta\overline{P}^{(k)},\overline{v}^{(k)}+\delta\overline{v}^{(k)},\overline{q}^{(k)}+\delta\overline{q}^{(k)}$;
  Update $\overline{\lambda}^{(k+1)}$ based on orthogonality by Assumption 1;
  Increment the iteration index $k\leftarrow k+1$;
@@
return Profile $\overline{U}^{\text{cv}}=\overline{U}^{(k)}$ and eigenpair $(\overline{\lambda},\overline{v}^{\text{cv}})=(\overline{\lambda}^{(k)},\overline{v}^{(k)})$;